# Sparse identification of nonlinear dynamics and Koopman operators with Shallow Recurrent Decoder Networks

### v1

@@ B.3 Flow over a cylinder
In the flow over a cylinder experiment, we follow the same settings as in the prior experiments and select the latent dimension to be 4. The forward integration time step is set to $dt=\frac{1}{300}$ corresponding to the frame rate of 30 FPS. We set the batch size at $64$ and the learning rate to $5e^{-4}$. The thresholding procedure is executed every 300 epochs with thresholds ranging from $(1e^{-4},1e^{-3})$.
B.4 Baseline experiment on pendulum
@@ B.4.1 Baseline methods and SINDy-SHRED setting
Appendix C Experiment on the 2D Kolmogorov flow
@@
In Eqn. 21, we find that $z_{1},z_{3},z_{5}$ are essentially a linear system. $z_{2},z_{4},z_{6}$ capture higher-order effects that are difficult to model without signal separation. We generate the trajectory from the initial condition at time point 0 and perform forward integration in Fig. 14. As we increase the Reynolds number, the discovery fails to produce robust predictions.
@@
Appendix D Sensor level plots of experiments
D.1 Sea surface temperature
@@ D.2 Ozone data
D.3 Flow over a cylinder
@@
D.4 Pendulum
@@ D.5 Kolmogorov flow
Appendix E Analysis of learned ODEs
E.1 Sea-surface temperature
The analytic solution to 11
@@ E.2 Ozone data
E.3 Flow over a cylinder

### v2

@@ D.2 Flow over a cylinder
In the flow over a cylinder experiment, we follow the same settings as in the prior experiments and select the latent dimension to be 4. The forward integration time step is set to $dt=\frac{1}{300}$ corresponding to the frame rate of 30 FPS. We set the batch size at $64$ and the learning rate to $5e^{-4}$. The thresholding procedure is executed every 300 epochs with thresholds ranging from $(1e^{-4},1e^{-3})$. From this extreme low-data limit, we manually perform data augmentation by reusing the latter part of the video once in the training loop to increase the number of available samples.
D.3 Baseline experiment on pendulum
@@
D.4 Sea-surface temperature data
Figure 10: Diagram of the RNN form of SINDy.
D.5 3D atmospheric ozone concentration
D.6 Isotropic turbulent flow
In the isotropic flow experiment, due to its complex nature, we select the latent dimension to be 16 with lag 100. The forward integration time step is set to $dt=0.0002$. We set the batch size to $128$ and the learning rate to $5e^{-4}$. The thresholding procedure is executed every 100 epochs with thresholds ranging from $(0.15,1.5)$. The total training epoch is 300.
Appendix E Experiment on the 2D Kolmogorov flow
@@
In Eqn. 72, we find that $z_{1},z_{3},z_{5}$ are essentially a linear system. $z_{2},z_{4},z_{6}$ capture higher-order effects that are difficult to model without signal separation. We generate the trajectory from the initial condition at time point 0 and perform forward integration in Fig. 11. As we increase the Reynolds number, the discovery fails to produce robust predictions.
@@
Appendix F Sensor level plots of experiments
F.1 Sea surface temperature
@@ F.2 Ozone data
F.3 Flow over a cylinder
Latent space of Koopman-SHRED
Figure 19: Extrapolation of latent representation in Koopman-SHRED from the discovered dynamical system for flow over a cylinder data. Colored: true latent representation. Grey: SINDy extrapolation.
@@ F.3 Flow over a cylinder
F.4 Isotropic turbulent flow
Figure 22: Extrapolation of all 16 latent representation in SINDy-SHRED from the discovered dynamical system for isotropic turbulent flow data for 250 frames. Colored: true latent representation. Grey: SINDy extrapolation.
Sensor-level prediction on the moving pendulum dataset.
Figure 23: Extrapolation of SINDy-SHRED for sensor-level predictions on the isotropic turbulent flow data. We randomly picked 18 sensors from spatial locations that are not in the sparse sensor training. The extrapolation shows the SINDy-SHRED prediction for the following 250 frames.
F.5 Pendulum
@@ F.6 Kolmogorov flow
Appendix G Analysis of learned ODEs
G.1 Sea-surface temperature
The analytic solution to 18
@@ G.2 Ozone data
G.3 Flow over a cylinder
@@
G.4 Isotropic turbulence flow
The complete closed form solution to Eq. equation 21 is given by:
$\mathbf{z}(t)=T\text{diag}\begin{pmatrix}e^{(0.47+9.39i)t}\\ e^{(0.47+9.39i)t}\\ e^{(0.05+11.90i)t}\\ e^{(0.05+11.90i)t}\\ e^{(0.03+13.42i)t}\\ e^{(0.03+13.42i)t}\\ e^{(-0.04+3.46i)t}\\ e^{(-0.04+3.46i)t}\\ e^{-0.26t}\\ e^{(-0.27+5.44i)t}\\ e^{(-0.27+5.44i)t}\\ e^{(-0.75+8.30i)t}\\ e^{(-0.75+8.30i)t}\\ e^{(-1.37+18.72i)t}\\ e^{(-1.37+18.72i)t}\\ e^{-3.39t}\\ \end{pmatrix}T^{-1}\mathbf{z}_{0}$ (76)
where $\mathbf{z}_{0}$ is the state at $t=0$ and $T$ contains all eigenvectors.